# Considerations for the planning, conduct and reporting of clinical trials with interim analyses

### v2

@@ Abstract
Interim analyses are prevalent in clinical trials. Although methodology is well established, there are aspects of how to operationalize and interpret interim analyses which remain unclear to many stakeholders. In this paper, a team of statisticians from the pharmaceutical industry, academia, and regulatory agencies provide a multi-stakeholder perspective on the key concepts behind interim analyses and considerations on terminology. We illustrate our proposals using a hypothetical clinical trial.
@@ 1 Introduction
Adaptive designs and especially interim analyses are prevalent in clinical trials [2]. Interim analyses are defined as pre-specified analyses that are performed during a clinical trial to determine whether criteria are met to implement pre-planned adjustments to the trial design [28, 8, 18]. Often, such criteria refer to potentially reaching an early conclusion on a scientific hypothesis. Different roles for interim analyses can be considered: (1) Interim analyses for efficacy, (2) interim analyses for futility, (3) interim analyses for safety and (4) interim analyses in adaptive designs where trial design features can be adapted, e.g., an arm is dropped, the sample size of the trial is modified, or one of several pre-specified populations is selected for enrichment. At a given interim analysis one or more of these assessments can be made.
@@
An interim analysis has implications for the conduct, reporting, and interpretation of a clinical trial. In our experience working as statisticians at regulatory agencies, in pharmaceutical companies, or in academia, we have observed that the design and statistical analysis plans of trials with interim analyses are at times poorly outlined and therefore the trials are inadequately reported. This makes trial designs and their results prone to misinterpretation within the trial team and by relevant external stakeholders such as regulatory agencies, HTA bodies, health care providers, investigators, patients and potentially even the wider public: An example of such misconceptions and misinterpretations based on the first four COVID-19 vaccines licensed in the EU is described by Hofner et al. [15].
This paper targets all stakeholders of clinical trials with interim analyses, i.e., statisticians and clinicians in pharmaceutical industry, regulatory agencies including HTA bodies, and academia. It is essential that all parties involved in planning, conduct, and interpretation of a clinical trial fully understand all the aspects related to the interim analysis strategy and that these are communicated in a transparent and precise manner to internal and external stakeholders. Currently, our impression is that there is a lack of standardized terminology which complicates the interpretation and clear communication of clinical trial results. Furthermore, the possible outcomes of interim analyses and their impact on trial conduct are often not clearly outlined. Through this paper, we recommend standard terminology and clarify aspects related to trial conduct and potential trial outcomes. We hope our paper facilitates: 1) informed decision making at the design stage; 2) trial conduct so that all involved parties have a full understanding of the intended scope of analyses; and 3) alignment and therefore a common interpretation of results across internal and external stakeholders.
Section 2 introduces a hypothetical trial that we then use in Section 3 to illustrate our proposed general terminology, while additional aspects are outlined in Section 4. We provide an example of text describing a trial design that could appear in a protocol or statistical analysis plan in Section 5. We illustrate reporting of such a trial in Section 6, and conclude with a discussion in Section 7.
@@ 2 Motivating Example – a hypothetical trial
We consider a hypothetical randomized controlled trial (RCT) with two treatment arms, experimental (E) vs. control (C). Let the primary endpoint be a time-to-event outcome, such as progression-free survival (PFS) or overall survival (OS) in oncology, or time to first relapse in multiple sclerosis. We will comment on more general endpoint definitions, multi-armed situations, and subgroup analyses in Section 7.
@@
The term minimal detectable difference refers to the minimal treatment effect which can still become statistically significant with the given sample size, or in other words the critical value of the hypothesis test on the scale of interest, here the hazard ratio. It may refer to an actual difference, e.g. between means, or for a relative risk measures such as the hazard ratio. To avoid statistically significant results that are clinically irrelevant it is of utmost importance to discuss with trial teams whether the minimal detectable difference is clinically relevant (0.8177 in our example), not necessarily whether the effect we power at (0.75 in our example) is clinically relevant. This has also been emphasized precisely in the context of a time-to-event endpoint in drug development by Carroll[4], where also formulas for minimal detectable difference computation for a time-to-event endpoint are provided. For the minimal detectable difference in the context of mean differences see e.g. Brock et al.[3].
To reduce the expected number of events (and hence to reduce the expected time needed to declare statistical significance) in case the experimental treatment does indeed work, we plan an interim analysis for efficacy at $2/3$ of the targeted number of events events. We use a Lan-DeMets $\alpha$-spending approximation [21] to O’Brien-Fleming [23] boundaries to control the overall type I error. To maintain the overall power, the maximum number of events is increased to 385. Consequently, with this interim analysis planned at an information fraction of $2/3$, it should be conducted after 257 of 385 events.
Similarly, to reduce the expected number of events if the intervention is not efficacious, we add a non-binding interim analysis for futility (see discussion in Section 3.5), the first after an information fraction of 1/3 (IA1, after 129 events) and the second after an information fraction of 2/3 (IA2, after 257 events), i.e., at the same time as the efficacy interim analysis. An independent data monitoring committee (iDMC, see below) is instructed to recommend stopping the trial for futility based on the observed hazard ratio if:
  • at IA1 the observed hazard ratio is $\geq 1$ and
@@
Assuming these decision criteria are followed, adding the futility analyses reduces the overall power of the trial from 80 to 78%. This is a direct consequence of the fact that even if the assumption on the alternative hypothesis is true, there are cases – albeit rather unlikely – were at the interim analyses the futility criterion would be met. It would be straightforward to compensate for this power loss by increasing the targeted number of events. For simplicity, and as the loss of power is minimal in that case, we have not done so here and our primary analysis remains planned at 385 events. Section 5 provides some more details and illustrates how we suggest to specify these interim analyses in a trial protocol or statistical analysis plan.
To ensure smooth conduct of the trial, the interim analyses should be carefully pre-planned and implications of interim decisions should be understood by all parties involved. In order to maintain trial integrity, interim analyses in registrational trials are typically conducted by an iDMC [18], which will provide recommendations to the sponsor of the trial. Further guidance on the installation and role of iDMCs are available from the FDA [27] and the EMA [7, 9], see also Schöffski[25].
General terminology for group-sequential designs, which will be helpful to align on at the design stage of the trial, is provided in Tables 1 and 2. Many of the terms in these tables (e.g., first-patient-in, clinical cutoff date CCOD, and snapshot data SSD) are also applicable to trials with other types of endpoints and without an interim analysis. Note that CCOD and SSD are analysis-specific. For each analysis, a corresponding CCOD and SSD is defined and the corresponding dataset is only generated at the SSD.
Notably, the simple term sample size may be confusing in communications across stakeholders for different endpoint definitions. The target sample size in a group-sequential design with a binary or continuous endpoint is the maximum number of patients. In trials with time-to-event endpoints, it is recommended to specifically use the terms number of events and number of patients and in fact an interim analysis does not usually affect the number of patients in the trial. The term sample size is a frequent source for confusion between the former and the latter.
@@
Group-sequential design A group-sequential design is a special adaptive design which allows for one or more interim analyses where the trial can potentially be stopped for efficacy, with rejection of the null hypothesis, or for futility, when the null hypothesis is unlikely to be rejected if the trial continues. See Table 5 for the design details of our hypothetical trial.
@@
Target information A group-sequential design is designed assuming a maximum number of patients (binary and continuous endpoint) or a maximum number of events (time-to-event endpoint) and this defines the CCOD of the primary analysis (see Tables 3 and 4 for a definition). This maximum number is target information and can be referred to also as the target number of patients or target number of events. Our hypothetical trial has a target number of events of 385, see Table 5.
Information fraction The proportion of patients or events available at an interim analysis relative to the target number of patients or events. In group-sequential designs, $\alpha$-spending functions generally specify the cumulative significance level at an interim analysis as a function of the information fraction. See Column Information Fraction in Table 5.
Independent data monitoring committee An iDMC is set up to ensure the integrity of the clinical trial. The iDMC has access to potentially unlinded data and gives recommendations based on pre-defined rules. Even in an open label trial a firewall should be in place to avoid access to data for the trial team.
@@
3 Types of Analyses
Moving through design, implementation and finally reporting of group-sequential designs, confusion can be generated through inappropriate use of seemingly “simple” standard terms, such as interim analysis, primary analysis or final analysis. In this section, we discuss some potential sources for confusion and propose alignment on terminology. Tables 3 and 4 define terms that relate to different analyses in a group-sequential design. Similar terms were previously defined by Hofner et al. [15] with the aim of demonstrating that standardized terminology would have enabled better informed discussions about the COVID-19 vaccine trials amongst all stakeholders. The first important distinction to draw is between the purpose of an analysis as envisaged at the design phase, and the actual purpose it has at the reporting stage of a trial. Furthermore, some analyses pre-specified during the design phase might not even take place because the trial is stopped early for either efficacy or futility. We discuss below what “stopping a trial” actually entails. Further analyses might be introduced to inform regulators or other stakeholders such as HTA bodies or prescribers and patients, either pre-planned or not. Hence, analyses may have separate but closely linked goals:
  • Hypothesis testing: Clinical trials are typically designed around the Neyman-Pearson hypothesis testing framework [22] with specification of a null and alternative hypothesis. Typically, formal trial success is based on hypothesis testing.
  • Estimation: Concerned with estimating an underlying population parameter and quantifying the uncertainty with which the trialist is able to do this.
While typically, a statistically significant hypothesis test can be considered as a minimum requirement for submission of a drug for regulatory approval, it is understood that in the comprehensive benefit-risk assessment precise estimation of the benefits and risks of the new medicine plays an equally crucial part. The regulatory approval of a drug is never a decision simply mapping the hypothesis test.
Care should be taken at the planning and reporting stages to clearly explain the value and intended scope of results, explain the distinction between hypothesis testing and estimation, and to avoid back-references to the analysis terminology used at the design stage.
@@ 3.2 Confirmatory analysis
Within a hypothesis test, the null hypothesis can only be rejected once for a given endpoint. Assume the null hypothesis for an endpoint is rejected at an interim analysis within a group-sequential design. We propose to refer to this analysis as confirmatory in the trial report and publications, instead of referring to it as interim analysis. The term interim analysis is potentially misleading as it suggests incomplete information and that a subsequent final analysis might overwrite the hypothesis decision. If the hypothesis was rejected at the planned primary analysis we propose to call this the confirmatory analysis as well. If the hypothesis was not rejected at any of the pre-specified interim analyses nor the primary analysis, the primary analysis is the confirmatory analysis which, in effect, confirmed the lack of a statistically significant outcome for the trial. It is of paramount importance to plan the group-sequential design such that the information available at any interim analysis and hence at each potential confirmatory analysis is sufficient for the intended purposes, e.g., regulatory submission. In particular, sufficient safety data and sufficiently mature data on other important efficacy endpoints (e.g., OS if PFS was the primary endpoint) should be available to support the risk-dimension of a benefit-risk assessment.
@@ 3.4 Decisive Analysis
3.5 Stopping a trial
In discussing trial designs across stakeholders, the formulation stopping for efficacy and stopping for futility is often used. Indeed, if the null hypothesis corresponding to the primary endpoint can be rejected at an interim analysis for efficacy within a group-sequential design, then the testing for this hypothesis is completed. In statistical terms, the trial is “stopped”. The result is statistically significant and the trial is formally a “success”. Similarily, if the trial is declared as futile, no further formal hypothesis testing will be conducted and, the result is “negative”.
From an operational perspective, however, the hypothesis test of the primary endpoint is only one piece of a clinical trial. Admittedly, given that it typically defines the design and “success” of a trial, it is an important one. However, data collection in the trial will typically continue after stopping for efficacy, i.e., rejection of the null hypothesis. This is done in order to refine estimation of the effect on the primary endpoint, and/or to collect more data to comprehensively describe and quantify the treatment effect, e.g., on secondary or safety endpoints (which might include formal hypothesis testing for these). An example is in oncology, where PFS may be the primary and OS the secondary endpoint. In some cases, a trial simply continues to allow patients to still receive treatment or patients are rolled over to an open-label extension study[5].
Stopping a trial for efficacy usually impacts the conduct of a trial even if the trial is not stopped from an operational perspective and further data is collected: Enrollment of new patients into the trial is usually stopped; Unblinded summary measures and results might be released; The trial sponsor might be unblinded to individual patient data; Investigators might be unblinded; Patients might be unblinded and/or be allowed to cross-over to the new therapy. All this has the potential to introduce bias and impact the continued collection of reliable and relevant data. Hence, care should be taken at the planning stage and measures should be considered to mitigate risks if possible or weigh them against potential benefits. These measures could include clear communication strategies, firewall procedures to protect the core trial team, and strategies how to continue the trial.
As discussed in Section 2, a futility analysis reduces the power of a hypothesis test, and therefore also the probability of a type 1 error. Hence, one might be tempted to adjust the critical value of the hypothesis test such that the significance level is again fully exploited, implying a slight power increase. In that case it is imperative to stick to the futility decision as otherwise the type 1 error would no longer be controlled. As the decision to stop a trial for futility is usually multi-faceted and includes amongst others also a review of the safety data, non-binding futility analyses are generally preferable[12].
@@ 3.5 Stopping a trial
Updated analysis Any analysis after the confirmatory analysis. It is recommended to pre-specify such analyses (including their timing and purpose) at the design stage in the protocol. Any analysis after the confirmatory analysis for a given endpoint that includes more data. This can include a pre-specified original interim or primary analysis (design stage), but also other analysis time points. It should be explicitly stated what analyses will be used as updated analyses. This may have to be defined conditionally on the time (interim analysis or primary analysis) when the null hypothesis for that endpoint was rejected, see Table 5.
@@
4 Subtleties
While the terminology above is relatively simple for a group-sequential design with a single hypothesis of interest, additional complexities arise in many situations, which will require further clarity in communication.

### v3

# Clinical trials with interim analyses: Standardizing Terminology to increase clarity

@@ Abstract
Interim analyses for group-sequential decision making are prevalent in clinical trials. Methodology is well established and has been routinely implemented over the last decades. Still, confusions and uncertainties on aspects of how to operationalize and interpret interim analyses exist for many stakeholders. In this paper, a team of statisticians from the pharmaceutical industry, academia, and regulatory agencies provide a multi-stakeholder perspective on the key concepts behind interim analyses, with the aim to introduce standard terminology to mitigate misunderstandings and facilitate clearer discussions.
@@ 1 Introduction
During the past decades, interim analyses implemented through a group-sequential or adaptive design have become common [2], especially for trials with time-to-event endpoints. The statistical theory is well established and straightforward to implement at the design stage. Terminology for such designs is, however, not standard, and to some extent, also misleading. Here, as a team of statisticians from pharmaceutical industry, regulatory agencies, and academia, we propose terminology for clinical trials with interim analyses and confirmatory intent, i.e. hypothesis testing. Interim analyses are defined as pre-specified analyses that are performed during a clinical trial to determine whether criteria are met to implement pre-planned adjustments to the trial design [30, 9, 21]. Often, such criteria refer to potentially reaching an early conclusion on a scientific hypothesis. Different roles for interim analyses can be considered: (1) Interim analyses for efficacy, (2) interim analyses for futility, (3) interim analyses for safety and (4) interim analyses in adaptive designs where trial design features can be adapted, e.g., an arm is dropped, the sample size of the trial is modified, or one of several pre-specified populations is selected for enrichment. At a given interim analysis one or more of these assessments can be made.
@@
An interim analysis has implications for the conduct, reporting, and interpretation of a clinical trial. In our experience working as statisticians at regulatory agencies, in pharmaceutical companies, or in academia, we have observed that the design and statistical analysis plans of trials with interim analyses are at times not appropriately outlined and therefore the trials are inadequately reported. Guidance on how to write statistical analysis plans exists, see e.g. Gamble et al.[15]. However, these guidances are not always followed and also do not require all the aspects that we propose to introduce for clearer communication. All this makes trial designs and their results prone to misinterpretation within the trial team and by relevant external stakeholders such as regulatory agencies, HTA (Health Technology Assessment) bodies, health care providers, investigators, patients and potentially even the wider public: an example of such misconceptions and misinterpretations based on the first four COVID-19 vaccines licensed in the EU is described by Hofner et al.[18]. This paper targets all stakeholders of clinical trials with interim analyses, i.e., statisticians and clinicians in the pharmaceutical industry, regulatory agencies including HTA bodies, and academia. It is essential that all parties involved in planning, conduct, and interpretation of a clinical trial fully understand all the aspects related to the interim analysis strategy and that these are communicated in a transparent and precise manner to internal and external stakeholders. Currently, our impression is that there is a lack of standardized terminology, including across pertinent (Health Authority) guidelines, which complicates the interpretation and clear communication of clinical trial results. Furthermore, the possible outcomes of interim analyses and their impact on trial conduct are often not clearly outlined. Through this paper, we recommend standard terminology and clarify aspects related to trial conduct and potential trial outcomes. This paper will facilitate: 1) informed decision making at the design stage; 2) trial conduct so that all involved parties have a full understanding of the intended scope of analyses; and 3) alignment and therefore a common interpretation of results across internal and external stakeholders. This work complements Ciolino et al (2023)[5] who provide a general overview and guidance on interim analyses for a nonstatistical audience. Based on our experience as various stakeholders involved in clinical trials at large, we focus on aspects of aligning terminology and clarifying communication of designs and results of clinical trials with interim analyses. As such, we target drug developers in general, including statisticians working on clinical trials. Section 2 introduces general concepts through a hypothetical trial that we then use in Section 3 to illustrate the core of the paper, our proposed general terminology, which we present in Tables 1 and 2. Additional aspects are outlined in Section 4. We conclude with a discussion in SectionA. We provide an example of text describing a trial design that could appear in a protocol or statistical analysis plan in an appendix to the paper. Also an example to illustrate reporting of such a trial is given in the appendix.
@@ 2 Motivating Example – a hypothetical trial
We consider a hypothetical randomized controlled trial (RCT) with two treatment arms, experimental (E) vs. control (C). Let the primary endpoint be a time-to-event outcome, such as progression-free survival (PFS) or overall survival (OS) in oncology, or time to first relapse in multiple sclerosis. We shall comment on more general endpoint definitions, multi-armed situations, and subgroup analyses in Section 5.
@@
The term minimal detectable difference refers to the minimal observed treatment effect which can still become statistically significant with the given sample size, or in other words the critical value of the hypothesis test on the scale of interest, here the hazard ratio. It may refer to an actual difference, e.g. between means, or to a relative risk measure such as the hazard ratio. To avoid statistically significant results that are not clinically relevant it is of importance to discuss with trial teams whether the minimal detectable difference is clinically relevant (0.8177 in our example), not necessarily whether the effect we power at (0.75 in our example) is clinically relevant. This has also been emphasized precisely in the context of a time-to-event endpoint in drug development by Carroll[4], where also formulas for minimal detectable difference computation for a time-to-event endpoint are provided. For the minimal detectable difference in the context of mean differences see, for example, Brock et al.[3].
To reduce the expected number of events (and hence to reduce the expected time needed to have the results available) in the case that the experimental treatment does indeed work, we plan an interim analysis for efficacy at $2/3$ of the targeted number of events events. We use a Lan-DeMets $\alpha$-spending approximation [24] to O’Brien-Fleming [25] boundaries to control the overall type I error. To maintain the overall power, the maximum number of events is increased to 385. Consequently, with this interim analysis planned at an information fraction of $2/3$, it should be conducted after 257 of 385 events.
Similarly, to reduce the expected number of events if the intervention is not efficacious, we add non-binding interim analyses for futility (see discussion in Section 4.3), the first after an information fraction of 1/3 (IA1, after 129 events) and the second after an information fraction of 2/3 (IA2, after 257 events), i.e., at the same time as the efficacy interim analysis. An independent data monitoring committee (iDMC, see below) is instructed to recommend stopping the trial for futility based on the observed hazard ratio if:
  • at IA1 the observed hazard ratio is $\geq 1$, or
@@
Assuming these decision criteria are followed, adding the futility analyses reduces the overall power of the trial from 80 to 78%. This is a direct consequence of the fact that even if the assumption on the alternative hypothesis is true, there are cases – albeit rather unlikely – where at the interim analyses the futility criterion would be met. It would be straightforward to compensate for this power loss by increasing the targeted number of events. For simplicity, and as the loss of power is minimal, the primary analysis remains planned at 385 events. Section A provides some more details and illustrates how we suggest to specify these interim analyses in a trial protocol or statistical analysis plan.
To ensure smooth conduct of the trial, the interim analyses should be carefully pre-planned and implications of interim decisions should be understood by all parties involved. In order to maintain trial integrity, interim analyses in registrational trials are typically conducted by an iDMC [21]. The iDMC will provide recommendations to the sponsor of the trial. Guidelines for those recommendations are typically described clearly in an iDMC-charter. If need be, the iDMC may deviate from the proposed recommendation guidelines, to recommend alternative actions (i.e. recommendation to add another interim analysis). Any recommendations from the iDMC are reduced to a minimum informative level (continue the trial unchanged, stop the trial for success) and are directed towards a sponsor committee, which itself is limited in size for integrity reasons. The sponsor may then follow the recommendations - or not. Further guidance on the installation and role of iDMCs are available from the FDA [29] and the EMA [8, 10], see also Schöffski[27]. Fleming et al.[12] provide additional guidance on best practices for iDMCs based on extensive experience.
General terminology for group-sequential designs, which will be helpful to align at the design stage of the trial, is provided in Tables 1 and 2. Many of the terms in these tables (e.g., first-patient-in, clinical cutoff date CCOD, and snapshot data SSD) are also applicable to trials with other types of endpoints and without an interim analysis. Note that CCOD and SSD are analysis-specific. For each analysis, a corresponding CCOD and SSD is defined and the corresponding dataset is only generated at the SSD.
Notably, the simple term sample size may be confusing in communications across stakeholders. The target sample size in a group-sequential design with a binary or continuous endpoint is the maximum number of patients. In trials with time-to-event endpoints, it is recommended to specifically use the terms number of events and number of patients and in fact an interim analysis does not usually affect the number of patients in the trial. The term sample size is a frequent source for confusion between the former and the latter.
@@
Group-sequential design (GSD) A group-sequential design is a special adaptive design which allows for one or more interim analyses where the trial can potentially be stopped for efficacy, with rejection of the null hypothesis, or for futility, when the null hypothesis is unlikely to be rejected if the trial continues. See Table 5 for the design details of our hypothetical trial.
@@
Target information GSDs are designed assuming a maximum number of patients (binary and continuous endpoint) or a maximum number of events (time-to-event endpoint). This maximum number is the target information (also referred to as target number of patients or target number of events). Our hypothetical trial has a target number of events of 385, see Table 5.
Information fraction The proportion of patients or events available at an interim analysis relative to the target number of patients or events. See Column Information Fraction in Table 5.
Independent data monitoring committee An iDMC is set up to ensure the integrity of the clinical trial. The iDMC has access to potentially unblinded data and gives recommendations based on pre-defined rules.
@@
3 Terminology for Different Types of Analyses
Moving through design, implementation and finally reporting of group-sequential designs, confusion can be generated through inappropriate use of seemingly “simple” standard terms, such as interim analysis, primary analysis or final analysis. In this section, we discuss some potential sources for confusion and propose alignment of terminology. Tables 3 and 4 define terms that relate to different analyses in group-sequential designs. Similar terms were previously defined by Hofner et al. [18] with the aim of demonstrating that standardized terminology would have enabled better informed discussions about the COVID-19 vaccine trials amongst all stakeholders. The first important distinction to draw is between the purpose of an analysis as envisaged at the design phase, and the actual purpose it has at the reporting stage of a trial. Furthermore, some analyses pre-specified during the design phase might not even take place because the trial is stopped early for either efficacy or futility. We discuss below what “stopping a trial” actually entails.
@@ 3.2 Confirmatory analysis
Within a hypothesis test, the null hypothesis can only be rejected once for a given endpoint. Assume the null hypothesis for an endpoint is rejected at an interim analysis within a group-sequential design. We propose to refer to this analysis as confirmatory in the trial report and publications, instead of referring to it as an interim analysis. The term interim analysis is potentially misleading as it suggests incomplete information and that a subsequent final analysis might overwrite the hypothesis decision. If the hypothesis was rejected at the planned primary analysis we propose to call this the confirmatory analysis as well. If the hypothesis was not rejected at any of the pre-specified interim analyses nor the primary analysis, the primary analysis is the confirmatory analysis which, in effect, confirmed the lack of a statistically significant outcome for the trial. It is of paramount importance to plan the group-sequential design such that the information available at any interim analysis and hence at each potential confirmatory analysis is sufficient for the intended purposes, e.g., regulatory submission. In particular, sufficient safety data and sufficiently mature data on other important efficacy endpoints (e.g., OS if PFS was the primary endpoint) should be available to support the risk-dimension of a benefit-risk assessment.
@@ 3.4 Decisive Analysis
Updated analysis Any analysis after the confirmatory analysis. It is recommended to pre-specify such analyses (including their timing and purpose) at the design stage in the protocol. Any analysis after the confirmatory analysis for a given endpoint that includes more data. This can include a pre-specified original interim or primary analysis (design stage), but also other analysis time points. It should be explicitly stated what analyses will be used as updated analyses. This may have to be defined conditionally on the time (interim analysis or primary analysis) when the null hypothesis for that endpoint was rejected, see Table A.
@@
4 Additional Clarifications on Group-Sequential Designs
The terminology above is relatively simple for a group-sequential design with a single hypothesis of interest. However, additional complexities arise in most situations in practice. These complexities may not be obvious in methodological developments which focus on single-endpoint situations, where all data is available as soon as a subject is randomized. In this section, we aim to clarify some additional subtle, yet still important, considerations to support development, discussion and implementation of well-designed group-sequential trials.